### v1

@@ Appendix F Interpretability
Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
Area TG dorsal Area PH $<0.001$ HC 0.684
Area TE1 anterior Area PH $<0.001$ HC 0.669
Area anterior 47r Area PH $<0.001$ HC 0.661
Area TG Ventral Area PH $<0.001$ HC 0.65
Primary Auditory Cortex Area $<0.001$ HC 0.646
Perirhinal Ectorhinal Cortex Area PH $<0.001$ HC 0.628
Primary Auditory Cortex Area V3CD $<0.001$ HC 0.625
Area TE2 anterior Area PH $<0.001$ HC 0.62
Orbital Frontal Complex Area PH $<0.001$ HC 0.614
Primary Auditory Cortex Area anterior 10p $<0.001$ HC 0.61
Primary Auditory Cortex Area PH $<0.001$ HC 0.602
Area TG dorsal Area V4t $<0.001$ HC 0.602
Dorsal Area 24d Area PH $<0.001$ HC 0.601
Area 10d Area PH $<0.001$ HC 0.594
ParaHippocampal Area 3 Area PH $<0.001$ HC 0.593
Area anterior 47r Area V4t $<0.001$ HC 0.592
Area Area PH $<0.001$ HC 0.585
ProStriate Area Area anterior 47r $<0.001$ HC 0.585
Area TE2 posterior Area V4t $<0.001$ HC 0.585
Area TG dorsal Area $<0.001$ HC 0.579
Area V4t Area PH $<0.001$ HC 0.579
Area 45 Area $<0.001$ HC 0.577
Area 47m Area $<0.001$ HC 0.577
Frontal Opercular Area 2 Area $<0.001$ HC 0.576
Primary Auditory Cortex Area anterior 47r $<0.001$ HC 0.575
Area TG Ventral Area V4t $<0.001$ HC 0.568
Polar 10p Area PH $<0.001$ HC 0.566
ProStriate Area Area V4t $<0.001$ HC 0.564
Area 47s Area PH $<0.001$ HC 0.564
Area TG dorsal Fusiform Face Complex $<0.001$ HC 0.563
Area anterior 47r Area $<0.001$ HC 0.558
Area STGa Area PH $<0.001$ HC 0.558
Area 10v Area PH $<0.001$ HC 0.557
Primary Auditory Cortex Area $<0.001$ HC 0.557
Frontal Opercular Area 3 Area $<0.001$ HC 0.553
Area OP1-SII Area anterior 47r $<0.001$ HC 0.552
Area TemporoParietoOccipital Junction 1 Area anterior 47r $<0.001$ HC 0.552
Superior Temporal Visual Area Area anterior 47r $<0.001$ HC 0.552
Area TE2 posterior Area PH $<0.001$ HC 0.55
ProStriate Area Area $<0.001$ HC 0.547
Fusiform Face Complex Area anterior 47r $<0.001$ HC 0.542
Area STGa Area $<0.001$ HC 0.541
Area posterior 47r Area PH $<0.001$ HC 0.54
Area 25 Area PH $<0.001$ HC 0.54
Auditory 4 Complex Area V3CD $<0.001$ HC 0.539
Area 10d Area V4t $<0.001$ HC 0.539
Primary Auditory Cortex Area V4t $<0.001$ HC 0.538
Area anterior 47r Area anterior 47r $<0.001$ HC 0.538
ParaHippocampal Area 3 Area anterior 47r $<0.001$ HC 0.536
Dorsal area 6 Area PH $<0.001$ HC 0.535
Area anterior 10p Area V4t $<0.001$ HC 0.535
Area 47m Area PH $<0.001$ HC 0.535
Area 8B Area PH $<0.001$ HC 0.535
Primary Visual Cortex Area PH $<0.001$ HC 0.534
Area STSv anterior Area PH $<0.001$ HC 0.534
Area TF Area PH $<0.001$ HC 0.534
Area V6A Area PH $<0.001$ HC 0.533
Area TG dorsal Area V3CD $<0.001$ HC 0.533
Middle Temporal Area Area PH $<0.001$ HC 0.533
Area OP2-3-VS Area anterior 47r $<0.001$ HC 0.533
Area 47s Area $<0.001$ HC 0.533
Posterior OFC Complex Area PH $<0.001$ HC 0.532
Perirhinal Ectorhinal Cortex Area V4t $<0.001$ HC 0.532
Area 10v Area $<0.001$ HC 0.532
Area posterior 10p Area PH $<0.001$ HC 0.531
Area anterior 10p Area PH $<0.001$ HC 0.531
Area 47l (47 lateral) Area PH $<0.001$ HC 0.53
Area 9 Middle Area PH $<0.001$ HC 0.53
Area STSd posterior Area PH $<0.001$ HC 0.53
Area TE2 posterior Area anterior 47r $<0.001$ HC 0.529
Area TE1 anterior Area V4t $<0.001$ HC 0.529
Area PHT Area PH $<0.001$ HC 0.528
Area PFt Area anterior 47r $<0.001$ HC 0.527
Area TE1 Middle Area PH $<0.001$ HC 0.526
Area STSd anterior Area PH $<0.001$ HC 0.526
Area 47s Area $<0.001$ HC 0.525
RetroSplenial Complex Area PH $<0.001$ HC 0.525
ParaHippocampal Area 1 Area PH $<0.001$ HC 0.525
Entorhinal Cortex Area PH $<0.001$ HC 0.524
Supplementary and Cingulate Eye Field Area $<0.001$ HC 0.524
Area PGp Area PH $<0.001$ HC 0.524
Area TG dorsal Area $<0.001$ HC 0.524
Dorsal area 6 Area $<0.001$ HC 0.524
Area anterior 47r Fourth Visual Area $<0.001$ HC 0.523
Area PFcm Area PH $<0.001$ HC 0.523
Supplementary and Cingulate Eye Field Area V3CD $<0.001$ HC 0.523
Area 10r Area PH $<0.001$ HC 0.522
Area Area PH $<0.001$ HC 0.521
Eighth Visual Area Area $<0.001$ HC 0.521
Area TE2 posterior Area $<0.001$ HC 0.52
VentroMedial Visual Area 1 Polar 10p $<0.001$ HC 0.519
Ventral Visual Complex Area TG Ventral $<0.001$ HC 0.519
Area 31p ventral Area PH $<0.001$ HC 0.519
Fourth Visual Area Area PH $<0.001$ HC 0.519
ParaHippocampal Area 3 Fourth Visual Area $<0.001$ HC 0.518
Area 7m Area PH $<0.001$ HC 0.518
Area p32 Area PH $<0.001$ HC 0.517
Auditory 4 Complex Area $<0.001$ HC 0.517
Area FST Area PH $<0.001$ HC 0.517
Area anterior 9-46v Area PH $<0.001$ HC 0.516
@@
Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
Area Lateral IntraParietal ventral VentroMedial Visual Area 3 $<0.001$ ADHD 1.109
Area s32 ParaHippocampal Area 3 $<0.001$ ADHD 1.062
Area TF Area 25 $<0.001$ ADHD 1.028
Inferior 6-8 Transitional Area VentroMedial Visual Area 3 $<0.001$ ADHD 1.016
Area TF Area s32 $<0.001$ ADHD 1.007
Ventral Visual Complex Area V3B $<0.001$ ADHD 1.002
Area TA2 VentroMedial Visual Area 3 $<0.001$ ADHD 1.002
Medial Area 7A VentroMedial Visual Area 3 $<0.001$ ADHD 1.0
Area 25 Area TF $<0.001$ ADHD 0.994
Area IntraParietal 0 VentroMedial Visual Area 3 $<0.001$ ADHD 0.994
Area 25 ParaHippocampal Area 3 $<0.001$ ADHD 0.993
Premotor Eye Field VentroMedial Visual Area 3 $<0.001$ ADHD 0.99
Area TF Orbital Frontal Complex $<0.001$ ADHD 0.99
Frontal Eye Fields VentroMedial Visual Area 3 $<0.001$ ADHD 0.979
Area 13l Area TF $<0.001$ ADHD 0.975
Area Lateral IntraParietal ventral Ventral Visual Complex $<0.001$ ADHD 0.969
Ventral Area 6 VentroMedial Visual Area 3 $<0.001$ ADHD 0.968
Frontal Eye Fields Eighth Visual Area $<0.001$ ADHD 0.968
Area 10r Area TF $<0.001$ ADHD 0.963
Medial Area 7A Eighth Visual Area $<0.001$ ADHD 0.96
Lateral Area 7P VentroMedial Visual Area 3 $<0.001$ ADHD 0.959
Inferior 6-8 Transitional Area Ventral Visual Complex $<0.001$ ADHD 0.955
Premotor Eye Field Ventral Visual Complex $<0.001$ ADHD 0.955
Area posterior 9-46v VentroMedial Visual Area 3 $<0.001$ ADHD 0.955
Medial IntraParietal Area VentroMedial Visual Area 3 $<0.001$ ADHD 0.954
Area 10r ParaHippocampal Area 3 $<0.001$ ADHD 0.953
Orbital Frontal Complex Area TF $<0.001$ ADHD 0.948
Frontal Eye Fields Posterior InferoTemporal complex $<0.001$ ADHD 0.948
Ventral IntraParietal Complex VentroMedial Visual Area 3 $<0.001$ ADHD 0.945
Area a24 Area TF $<0.001$ ADHD 0.944
Area IFSp VentroMedial Visual Area 3 $<0.001$ ADHD 0.939
Area 6 anterior VentroMedial Visual Area 3 $<0.001$ ADHD 0.938
Frontal Eye Fields Ventral Visual Complex $<0.001$ ADHD 0.934
Area 8BM Area TF $<0.001$ ADHD 0.934
Area IntraParietal 1 Area TF $<0.001$ ADHD 0.934
Area s32 Area TF $<0.001$ ADHD 0.934
Posterior InferoTemporal complex Orbital Frontal Complex $<0.001$ ADHD 0.933
Area V3B VentroMedial Visual Area 3 $<0.001$ ADHD 0.932
IntraParietal Sulcus Area 1 VentroMedial Visual Area 3 $<0.001$ ADHD 0.932
Area TF Area V3B $<0.001$ ADHD 0.932
Area IFJa Posterior InferoTemporal complex $<0.001$ ADHD 0.93
IntraParietal Sulcus Area 1 Ventral Visual Complex $<0.001$ ADHD 0.927
Area a24 ParaHippocampal Area 3 $<0.001$ ADHD 0.927
Area 6 anterior Ventral Visual Complex $<0.001$ ADHD 0.926
Area 46 Eighth Visual Area $<0.001$ ADHD 0.925
Area Lateral IntraParietal ventral Fourth Visual Area $<0.001$ ADHD 0.923
Inferior 6-8 Transitional Area ParaHippocampal Area 3 $<0.001$ ADHD 0.923
Area IFJp VentroMedial Visual Area 3 $<0.001$ ADHD 0.921
Lateral Area 7P Ventral Visual Complex $<0.001$ ADHD 0.92
Inferior 6-8 Transitional Area VentroMedial Visual Area 2 $<0.001$ ADHD 0.92
Area dorsal 32 Area TF $<0.001$ ADHD 0.92
Area TF Area a24 $<0.001$ ADHD 0.919
Area IntraParietal 1 VentroMedial Visual Area 3 $<0.001$ ADHD 0.918
@@
Posterior OFC Complex Area TF $<0.001$ ADHD 0.917
Inferior 6-8 Transitional Area Eighth Visual Area $<0.001$ ADHD 0.917
Area 10r Posterior InferoTemporal complex $<0.001$ ADHD 0.916
Area IFSp Posterior InferoTemporal complex $<0.001$ ADHD 0.916
Area 6 anterior Area Lateral Occipital 2 $<0.001$ ADHD 0.916
Anterior 24 prime VentroMedial Visual Area 3 $<0.001$ ADHD 0.913
Area IFJa Area TF $<0.001$ ADHD 0.912
Area s32 Posterior InferoTemporal complex $<0.001$ ADHD 0.911
Area IntraParietal 0 ParaHippocampal Area 3 $<0.001$ ADHD 0.91
Area 55b VentroMedial Visual Area 3 $<0.001$ ADHD 0.907
Inferior 6-8 Transitional Area Fourth Visual Area $<0.001$ ADHD 0.907
Anterior Ventral Insular Area Posterior InferoTemporal complex $<0.001$ ADHD 0.906
Area Lateral IntraParietal ventral Eighth Visual Area $<0.001$ ADHD 0.906
Area p32 prime VentroMedial Visual Area 3 $<0.001$ ADHD 0.905
Medial IntraParietal Area Ventral Visual Complex $<0.001$ ADHD 0.904
Area Lateral IntraParietal ventral VentroMedial Visual Area 1 $<0.001$ ADHD 0.904
Area p32 Area TF $<0.001$ ADHD 0.903
Area TA2 Eighth Visual Area $<0.001$ ADHD 0.902
Area 8Ad Area TF $<0.001$ ADHD 0.902
Area Lateral IntraParietal ventral VentroMedial Visual Area 2 $<0.001$ ADHD 0.902
Area 6 anterior Eighth Visual Area $<0.001$ ADHD 0.901
Premotor Eye Field Eighth Visual Area $<0.001$ ADHD 0.899
Area STSv anterior Area TF $<0.001$ ADHD 0.899
Medial Area 7A Ventral Visual Complex $<0.001$ ADHD 0.897
Posterior OFC Complex Posterior InferoTemporal complex $<0.001$ ADHD 0.896
Area 55b Posterior InferoTemporal complex $<0.001$ ADHD 0.895
Area 25 Posterior InferoTemporal complex $<0.001$ ADHD 0.894
Area PF Opercular VentroMedial Visual Area 3 $<0.001$ ADHD 0.893
Area IFJp VentroMedial Visual Area 1 $<0.001$ ADHD 0.891
Ventral IntraParietal Complex Ventral Visual Complex $<0.001$ ADHD 0.891
Anterior Agranular Insula Complex VentroMedial Visual Area 3 $<0.001$ ADHD 0.891
Area 55b Ventral Visual Complex $<0.001$ ADHD 0.89
Entorhinal Cortex ParaHippocampal Area 3 $<0.001$ ADHD 0.889
Area 46 VentroMedial Visual Area 3 $<0.001$ ADHD 0.889
Area STSd anterior Area TF $<0.001$ ADHD 0.887
Area TF Area IntraParietal 2 $<0.001$ ADHD 0.887
Area 6 anterior ParaHippocampal Area 3 $<0.001$ ADHD 0.887
Area IFJp Posterior InferoTemporal complex $<0.001$ ADHD 0.886
Frontal Opercular Area 4 VentroMedial Visual Area 3 $<0.001$ ADHD 0.886
Area IFSp Ventral Visual Complex $<0.001$ ADHD 0.884
Middle Insular Area VentroMedial Visual Area 3 $<0.001$ ADHD 0.883
Medial IntraParietal Area Area TF $<0.001$ ADHD 0.881
Ventral Visual Complex Frontal Eye Fields $<0.001$ ADHD 0.881
Area 7m Area TF $<0.001$ ADHD 0.879
Area 6m anterior Eighth Visual Area $<0.001$ ADHD 0.879
Area Lateral IntraParietal dorsal Ventral Visual Complex $<0.001$ ADHD 0.879
Area TA2 Posterior InferoTemporal complex $<0.001$ ADHD 0.878

### v2

@@ Appendix F Interpretability
Region 1 Region 2 Disorder/HC Corrected p-value Cohen’s d
PeriSylvian Language Area L Area OP2-3-VS R HC $<0.001$ 0.441
Dorsal Area 24d L Area OP2-3-VS R HC $<0.001$ 0.441
Area 31pd L Auditory 4 Complex R HC $<0.001$ 0.44
Area STSv posterior R Auditory 4 Complex R HC $<0.001$ 0.44
Area STSv posterior R Area OP2-3-VS R HC $<0.001$ 0.439
Dorsal area 6 R Area OP2-3-VS R HC $<0.001$ 0.435
Area TE1 Middle L Frontal Opercular Area 2 R HC $<0.001$ 0.435
Area TE1 anterior R Auditory 4 Complex R HC $<0.001$ 0.434
Area STSd anterior L Area OP2-3-VS R HC $<0.001$ 0.433
Auditory 5 Complex R Area OP2-3-VS R HC $<0.001$ 0.431
Superior Temporal Visual Area L Insular Granular Complex L HC $<0.001$ 0.43
RetroInsular Cortex L Area OP2-3-VS R HC $<0.001$ 0.429
Area TE1 anterior R Area OP2-3-VS R HC $<0.001$ 0.429
Area STSd posterior L Area OP2-3-VS R HC $<0.001$ 0.428
Area 23d L Frontal Opercular Area 2 R HC $<0.001$ 0.424
Area TG dorsal L Insular Granular Complex L HC $<0.001$ 0.423
Sixth Visual Area R Area OP2-3-VS R HC $<0.001$ 0.423
PeriSylvian Language Area L Insular Granular Complex L HC $<0.001$ 0.423
PeriSylvian Language Area L Frontal Opercular Area 2 R HC $<0.001$ 0.422
Entorhinal Cortex R Area OP2-3-VS R HC $<0.001$ 0.422
PeriSylvian Language Area L Auditory 4 Complex R HC $<0.001$ 0.422
Area STSv anterior L Area OP2-3-VS R HC $<0.001$ 0.421
Area s32 L Frontal Opercular Area 2 R HC $<0.001$ 0.421
Area s32 L Area OP2-3-VS R HC $<0.001$ 0.421
Area TE1 Middle L Area OP2-3-VS R HC $<0.001$ 0.42
Area 8Av L Insular Granular Complex L HC $<0.001$ 0.42
Area STSd anterior R Area OP2-3-VS R HC $<0.001$ 0.418
Area PHT R Area OP2-3-VS R HC $<0.001$ 0.417
VentroMedial Visual Area 3 L Insular Granular Complex L HC $<0.001$ 0.417
Area V6A L Area OP2-3-VS R HC $<0.001$ 0.417
Area 47l (47 lateral) R Area OP2-3-VS R HC $<0.001$ 0.416
Dorsal area 6 L Area OP2-3-VS R HC $<0.001$ 0.415
Area 8Ad L Insular Granular Complex L HC $<0.001$ 0.414
Area 8C L Insular Granular Complex L HC $<0.001$ 0.414
Area TE1 anterior L Auditory 4 Complex R HC $<0.001$ 0.413
Area V6A L Frontal Opercular Area 2 R HC $<0.001$ 0.413
Area TG dorsal L Auditory 4 Complex R HC $<0.001$ 0.413
Auditory 5 Complex L Area OP2-3-VS R HC $<0.001$ 0.412
Area 10d R Frontal Opercular Area 2 R HC $<0.001$ 0.412
Area 47s R Frontal Opercular Area 2 R HC $<0.001$ 0.411
Area TE1 Middle L Auditory 4 Complex R HC $<0.001$ 0.411
Area TE1 posterior L Insular Granular Complex L HC $<0.001$ 0.411
Hippocampus R Auditory 4 Complex R HC $<0.001$ 0.411
Area TE1 Middle L Insular Granular Complex L HC $<0.001$ 0.41
Area STSd posterior L Auditory 4 Complex R HC $<0.001$ 0.41
Primary Auditory Cortex R Auditory 4 Complex R HC $<0.001$ 0.41
Dorsal Area 24d L Auditory 4 Complex R HC $<0.001$ 0.41
Area dorsal 32 L Area OP2-3-VS R HC $<0.001$ 0.41
Area TE1 posterior R Area OP2-3-VS R HC $<0.001$ 0.409
Dorsal area 6 R Insular Granular Complex L HC $<0.001$ 0.409
Primary Motor Cortex R Auditory 4 Complex R HC $<0.001$ 0.409
Area TemporoParietoOccipital Junction 1 R Frontal Opercular Area 2 R HC $<0.001$ 0.409
Area TE1 anterior R Frontal Opercular Area 2 R HC $<0.001$ 0.409
Area TE1 Middle R Area OP2-3-VS R HC $<0.001$ 0.408
Dorsal area 6 R Frontal Opercular Area 2 R HC $<0.001$ 0.408
Lateral Belt Complex R Auditory 4 Complex R HC $<0.001$ 0.408
Area 31pd L Insular Granular Complex L HC $<0.001$ 0.408
Superior 6-8 Transitional Area L Insular Granular Complex L HC $<0.001$ 0.408
Lateral Belt Complex R Area OP2-3-VS R HC $<0.001$ 0.407
Area TG dorsal R Auditory 4 Complex R HC $<0.001$ 0.407
Area 31pd L Area OP2-3-VS R HC $<0.001$ 0.407
Area dorsal 32 L Frontal Opercular Area 2 R HC $<0.001$ 0.407
Area 10d R Area OP2-3-VS R HC $<0.001$ 0.407
Area STSd anterior L Insular Granular Complex L HC $<0.001$ 0.406
Area 8B Lateral L Insular Granular Complex L HC $<0.001$ 0.405
Area 31pd R Auditory 4 Complex R HC $<0.001$ 0.405
Area 8Ad L Area OP2-3-VS R HC $<0.001$ 0.405
Pirform Cortex L Area OP2-3-VS R HC $<0.001$ 0.405
Frontal Opercular Area 1 R Area OP2-3-VS R HC $<0.001$ 0.405
Area TemporoParietoOccipital Junction 1 R Auditory 4 Complex R HC $<0.001$ 0.405
Dorsal area 6 R Medial Belt Complex R HC $<0.001$ 0.405
Area 8B Lateral L Area OP2-3-VS R HC $<0.001$ 0.404
Area PGs L Area OP2-3-VS R HC $<0.001$ 0.404
Area TemporoParietoOccipital Junction 1 R Area OP2-3-VS R HC $<0.001$ 0.404
Area 6mp R Auditory 4 Complex R HC $<0.001$ 0.403
Area IFJa R Auditory 4 Complex R HC $<0.001$ 0.403
Area 45 L Insular Granular Complex L HC $<0.001$ 0.403
Area 31pd R Insular Granular Complex L HC $<0.001$ 0.403
Ventral IntraParietal Complex R Area OP2-3-VS R HC $<0.001$ 0.403
Auditory 5 Complex L Insular Granular Complex L HC $<0.001$ 0.403
Auditory 4 Complex L Area OP2-3-VS R HC $<0.001$ 0.403
Area PFcm R Auditory 4 Complex R HC $<0.001$ 0.403
Area 6 anterior L Area OP2-3-VS R HC $<0.001$ 0.403
Area dorsal 32 L Insular Granular Complex L HC $<0.001$ 0.403
Area 10d L Area OP2-3-VS R HC $<0.001$ 0.403
Area 9 Posterior L Insular Granular Complex L HC $<0.001$ 0.402
Middle Temporal Area R Area OP2-3-VS R HC $<0.001$ 0.402
Area 8Ad L Frontal Opercular Area 2 R HC $<0.001$ 0.402
Area TE1 Middle L Posterior Insular Area 2 R HC $<0.001$ 0.401
Area STSd anterior R Auditory 4 Complex R HC $<0.001$ 0.401
Area TG dorsal R Frontal Opercular Area 2 R HC $<0.001$ 0.401
Area TE1 posterior R Frontal Opercular Area 2 R HC $<0.001$ 0.401
Area 7m R Auditory 4 Complex R HC $<0.001$ 0.401
Entorhinal Cortex R Auditory 4 Complex R HC $<0.001$ 0.401
Area 47l (47 lateral) L Area OP2-3-VS R HC $<0.001$ 0.4
PreCuneus Visual Area L Insular Granular Complex L HC $<0.001$ 0.4
Area STGa R Area OP2-3-VS R HC $<0.001$ 0.4
Superior Temporal Visual Area L Area OP2-3-VS R HC $<0.001$ 0.4
Area TG dorsal L Area OP2-3-VS R HC $<0.001$ 0.4
Medial Belt Complex R Area OP2-3-VS R HC $<0.001$ 0.399
@@
Region 1 Region 2 Disorder/HC Corrected p-value Cohen’s d
Medial Belt Complex R Area PGp R HC $<0.001$ 0.373
Area dorsal 32 R ParaHippocampal Area 1 R ADHD $<0.001$ 0.372
Area FST R ParaHippocampal Area 1 R ADHD $<0.001$ 0.368
Area dorsal 32 R Area OP2-3-VS R ADHD $<0.001$ 0.367
Medial Belt Complex R Seventh Visual Area R HC $<0.001$ 0.363
Area PGi L ParaHippocampal Area 1 R ADHD $<0.001$ 0.361
ParaHippocampal Area 3 R Area TF R HC 0.001 0.346
Area 47s R Area PGp R HC 0.001 0.344
Area 8BM R Area OP2-3-VS R ADHD 0.001 0.343
Area FST R Area OP2-3-VS R ADHD 0.001 0.343
Area 11l L Area PGp R HC 0.001 0.343
Area 8Av R ParaHippocampal Area 1 R ADHD 0.001 0.343
Area 11l L Anterior IntraParietal Area L HC 0.001 0.342
Area 47l (47 lateral) L Anterior IntraParietal Area L HC 0.001 0.341
Area PGi L Area OP2-3-VS R ADHD 0.001 0.339
Area 8BM R ParaHippocampal Area 1 R ADHD 0.001 0.339
Area FST R Area Posterior Insular 1 L ADHD 0.001 0.338
Area TemporoParietoOccipital Junction 2 L ParaHippocampal Area 1 R ADHD 0.001 0.337
Area 47s R Anterior IntraParietal Area L HC 0.001 0.336
Area OP4-PV R Area PGp R HC 0.001 0.336
Pirform Cortex R Area PGp R HC 0.001 0.336
Dorsal Transitional Visual Area R ParaHippocampal Area 1 R ADHD 0.001 0.335
Area dorsal 23 a+b R ParaHippocampal Area 1 R ADHD 0.001 0.334
Fourth Visual Area L Area OP2-3-VS R ADHD 0.001 0.334
Area STSv posterior R Area 47s R ADHD 0.001 0.333
Area Posterior Insular 1 L Medial Superior Temporal Area L HC 0.001 0.33
Area 47s R Seventh Visual Area R HC 0.001 0.33
Area STSv posterior R Polar 10p R ADHD 0.001 0.33
Area 11l L Seventh Visual Area R HC 0.001 0.33
Area OP4-PV R Area TF R HC 0.001 0.329
Medial Belt Complex R Sixth Visual Area L HC 0.001 0.328
@@
Area STSv posterior R Anterior Agranular Insula Complex R ADHD 0.002 0.328
Area STSv posterior R ParaHippocampal Area 1 R ADHD 0.002 0.328
Area 5L R ParaHippocampal Area 1 R ADHD 0.002 0.328
Area IFSa R Area OP2-3-VS R ADHD 0.002 0.327
Medial Belt Complex R Medial Superior Temporal Area L HC 0.002 0.326
Area STGa R Area TF R HC 0.002 0.325
Area 11l L Area 7PC L HC 0.002 0.325
Area V3A L Area OP2-3-VS R ADHD 0.002 0.324
Area posterior 9-46v R Area OP2-3-VS R ADHD 0.002 0.324
Area anterior 9-46v R Area OP2-3-VS R ADHD 0.002 0.324
Fourth Visual Area R Area 10v L ADHD 0.002 0.323
Medial Belt Complex R Anterior IntraParietal Area L HC 0.002 0.323
Area STSv posterior R Orbital Frontal Complex L ADHD 0.002 0.323
Area PFm Complex L Area OP2-3-VS R ADHD 0.002 0.322
Area STSv posterior R Area OP2-3-VS R ADHD 0.002 0.322
Medial Area 7A R ParaHippocampal Area 1 R ADHD 0.002 0.321
Medial Belt Complex R Area V3A L HC 0.002 0.32
Area OP4-PV R Medial Superior Temporal Area L HC 0.002 0.32
Area STSv posterior R Area 10v L ADHD 0.002 0.32
Area 11l L Fourth Visual Area L HC 0.002 0.32
Area STSd posterior L ParaHippocampal Area 1 R ADHD 0.002 0.319
Area dorsal 32 R Area Posterior 24 prime R ADHD 0.002 0.319
Area dorsal 32 R Hippocampus R ADHD 0.002 0.319
Area Posterior Insular 1 L Anterior IntraParietal Area L HC 0.002 0.318
Area PFm Complex L ParaHippocampal Area 1 R ADHD 0.002 0.318
Third Visual Area R Area OP2-3-VS R ADHD 0.002 0.318
Area FST L ParaHippocampal Area 1 R ADHD 0.002 0.317
Area FST R Pirform Cortex L ADHD 0.002 0.317
Area FST R Area p32 L ADHD 0.002 0.317
Anterior IntraParietal Area L ParaHippocampal Area 1 R ADHD 0.002 0.316
Area 11l L Ventral IntraParietal Complex L HC 0.002 0.316
Area STSv posterior R Area 47l (47 lateral) R ADHD 0.002 0.316
Area STSd posterior L Area OP2-3-VS R ADHD 0.002 0.316
Area ventral 23 a+b R ParaHippocampal Area 1 R ADHD 0.002 0.315
Area STSv posterior R Pirform Cortex L ADHD 0.002 0.315
Medial Belt Complex R Second Visual Area L HC 0.002 0.315
Area Posterior Insular 1 L Area PGp R HC 0.002 0.315
Area 47l (47 lateral) L Medial Superior Temporal Area L HC 0.002 0.315
Area 8Av R Area OP2-3-VS R ADHD 0.002 0.314
Pirform Cortex R Seventh Visual Area R HC 0.002 0.314
Sixth Visual Area L ParaHippocampal Area 1 R ADHD 0.002 0.314
Area 43 L Area OP2-3-VS R ADHD 0.002 0.313
Area STSv posterior R Area p32 L ADHD 0.002 0.313
Area TG dorsal R Area PGp R HC 0.002 0.313
Medial Superior Temporal Area L ParaHippocampal Area 1 R ADHD 0.003 0.312
Area dorsal 32 R Anterior Agranular Insula Complex R ADHD 0.002 0.312
Area p32 R ParaHippocampal Area 1 R ADHD 0.003 0.312
Area PGi L Hippocampus R ADHD 0.003 0.312
Area STGa L Anterior IntraParietal Area L HC 0.002 0.312
Medial Area 7A R Area OP2-3-VS R ADHD 0.003 0.311
Dorsal area 6 R Medial Superior Temporal Area L HC 0.003 0.311
Area TE1 Middle R ParaHippocampal Area 1 R ADHD 0.003 0.311
Area p32 L Medial Superior Temporal Area L HC 0.003 0.31
IntraParietal Sulcus Area 1 R ParaHippocampal Area 1 R ADHD 0.003 0.31
Fusiform Face Complex R ParaHippocampal Area 1 R ADHD 0.003 0.309
Area 11l L Medial Superior Temporal Area L HC 0.003 0.309
Area 44 R ParaHippocampal Area 1 R ADHD 0.003 0.309
Area 23c R ParaHippocampal Area 1 R ADHD 0.003 0.309
Medial Belt Complex R Area V6A R HC 0.003 0.309
Parieto-Occipital Sulcus Area 1 R Anterior IntraParietal Area L HC 0.003 0.308
Medial Belt Complex R Ventral IntraParietal Complex L HC 0.003 0.308
Dorsal area 6 R Seventh Visual Area R HC 0.003 0.308
Area 47l (47 lateral) L Area PGp R HC 0.003 0.308
Area STSv posterior R Area anterior 10p L ADHD 0.003 0.307
Ventral Area 6 L Medial Superior Temporal Area L HC 0.003 0.306
Area TemporoParietoOccipital Junction 1 R ParaHippocampal Area 1 R ADHD 0.003 0.306
Pirform Cortex L Area PGp R HC 0.003 0.306
Area STSv posterior R Area 47s L ADHD 0.003 0.305
Parieto-Occipital Sulcus Area 2 L ParaHippocampal Area 1 R ADHD 0.003 0.305